OpenStudy (anonymous):

by selling a television set for $260 a dealer finds he is making a profit of 30% of cost. at what price must he sell it to make a profit of 40% of cost? A. $196 B. $254.80 C. $280 D. $282 E. $322

OpenStudy (anonymous):

x+x(.3)=260 solve for x

OpenStudy (anonymous):

then when you have x, use the same equation but put .4 in where .3 is and solve

OpenStudy (anonymous):

do u divide 260 by .3 first?

OpenStudy (anonymous):

no, add the x's first

OpenStudy (anonymous):

i got 2x(.3)=260 now what do i have to do?

OpenStudy (anonymous):

no, because it is x + .3x, you get 1.3x. then divide 260 by 1.3 which gives you 200.

OpenStudy (anonymous):

but 200 is not in the choices

OpenStudy (anonymous):

i know. you have to take 200 and fill it back into the original equation for x. but where we had .3, you now need .4

OpenStudy (anonymous):

260=x+x(.4)

OpenStudy (anonymous):

260=1.4x?

OpenStudy (anonymous):

no, x= 200+200(.4)

OpenStudy (anonymous):

you were close

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

you have to work backwards to find the original price, 200, then go forward again using 40 percent instead of 30 percent

OpenStudy (anonymous):

i got $280

OpenStudy (anonymous):

great job